the hypotenuse of a right triangle is 28cm long and the length of one of the legs is 23cm, find the length of the other leg

Answer by TimothyLamb(4379)

use pythagorean thm:
h = sqrt( aa + bb )
28 = sqrt( 23^2 + b^2 )
28^2 = 23^2 + b^2
784 = 529 + b^2
b^2 = 784 - 529
b^2 = 255
---
b = 15.97 cm
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php